# Infrared-Radio-follow-up Observations for Detection of the Magnetic Radio Emission of Extra-Solar Planets: A New Window to Detect Exoplanets

### v1

@@ Abstract
There are several methods for indirectly detecting exoplanets, such as transit, radial velocity, astrometry, and the conventional gravitational microlensing approach. These methods rely on observing the effects of exoplanets on the emission or motion of observed stars. All these techniques have focused on the optical or infrared domains. However, an alternative method for exoplanet detection via microlensing events involves planets orbiting the source star, creating a binary source system. In this study, we explore a novel approach to detecting and studying exoplanets exclusively through their radio emissions resulting from magnetospheric processes. We propose utilizing the Roman telescope as a survey observer to detect microlensing events. Subsequently, we investigate the potential for detecting planetary radio signals through follow-up observations of these microlensing events in the radio band using the SKA telescope. This method is viable due to the comparable radio emission levels of exoplanets and their parent stars, unlike optical and infrared emissions. We conduct a Monte Carlo simulation to replicate the observations by the Nancy Roman Telescope, followed by a follow-up observation in radio frequencies using the SKA telescope. We determine that approximately 1155 exoplanets exhibit detectable signals by the SKA telescope during the 7-season observations by the Nancy Roman Telescope. This result indicates that such a method cannot only facilitate the direct detection of exoplanets but also enable the measurement of their magnetic field strength through analysis of their radio emissions.
journal: Frontiers
@@ 1 Introduction
A planetary magnetic field is one of the few observable quantities that contain information about the properties of the planetary interior since the internal dynamo produces it. Further, the extent to which exoplanets retain their atmospheres over geological time scales depends on whether they are exposed directly to stellar winds. Therefore, information about the magnetic fields of extrasolar planets is essential to assess their habitability. So far, our knowledge of planetary magnetic fields is limited to our solar system’s planets, which are not representative of the vast array of possible planets, such as hot Jupiters.
@@
In this work, we combine two different exoplanet detection methods to make the radio emission of exoplanets detectable. The novel idea behind this study is that during a microlensing event, light intensity will be magnified at all frequencies. Therefore, a microlensing event can serve as a natural light magnifier. The idea is to use an optical/infrared telescope to detect a microlensing event that would trigger follow-up observations by radio telescopes upon detection. If the source star has at least one exoplanet in microlensing events, we can directly measure the light emitted from the exoplanet (1). The magnified light of the exoplanet could be observable in the infrared wavelength as we have shown in (Bagheri et al., 2019). However, the radio band is the most desirable wavelength window to detect exoplanets (Vedantham et al., 2020; Viswanath et al., 2020; Foster et al., 2020) since the planetary radio emission, which includes the effects of exomoons, may dominate the signal from the parent star (Cauley et al., 2019; Zarka, 2004). This method will enable us to measure the magnetic fields of exoplanets and provide us with information about the existence of exomoons. To show this, we simulate the Nancy Grace Roman Space Telescope (previously named Wide Filed Infrared Space Telescope, WFIRST) as an infrared survey telescope to alert the Square Kilometre Array (SKA) radio telescope to the occurrence of microlensing events for follow-up observations.
@@ 2.1 Microlensing events
where $u$ is the impact parameter, defined as the closest approach of the source to the lens normalized by the Einstein radius. The lensing system may comprise either a single object (single-lens) or two objects (binary-lens), with the possibility of additional components (Daněk & Heyrovskỳ, 2019). The presence of two lenses introduces additional complexity to the gravitational lensing phenomenon, leading to more intricate and often asymmetric light curves. Source systems can also consist of one or more objects. The conventional technique utilized in gravitational microlensing for exoplanet detection relies on the premise that the lens star hosts an exoplanet, constituting a binary lens system. Alternatively, another microlensing approach involves a scenario where the planet orbits the source star, resulting in a binary source configuration (Bagheri et al., 2019) as illustrated in 1. In this study we simulate binary-source micorlensing events by assuming both single-lens and binary-lens configurations. We consider the finite source effect for the source star while neglecting limb darkening.
@@ 2.3 Exoplanet Parameters and their Occurrence
In our simulations, we categorize the characteristic parameters of exoplanets into two primary groups: rocky planets and Jovian planets. According to the conventional planet formation theory, rocky planets are situated within the snow line of their parent star (Kennedy et al., 2006; Kennedy & Kenyon, 2008). This snowline can be expressed with the snowline of the Sun as
@@
The planetary mass range in our simulation is $0.002-13.5M_{J}$. For giant planets orbiting host stars with masses greater than $0.6$ solar masses, we utilized the correlation between giant planet occurrence and host star mass and metallicity as outlined by (Johnson et al., 2010). This relationship was established for stars within the range of $0.5$ to $2.0$ solar masses ($0.5M_{\odot}<M_{\star}<2.0M_{\odot}$) and metallicities below $0.4([Fe/H]<0.4)$, but (Perryman et al., 2014) extrapolated this relationship for masses greater than two solar mass as the occurrence rate of giant planets around stars with masses exceeding 2.0 solar masses ($M_{\star}>2.0M_{\odot}$) equals that at $M_{\star}=2.0M_{\odot}$, and likewise, the rate for stars with metallicities exceeding 0.4 ($[Fe/H]>0.4$) equals that at $[Fe/H]=0.4$. Following the occurrence dependencies discussed in (Perryman et al., 2014), we adopted their result of planet occurrence as summarized in table LABEL:tab:fvsMp.
@@
$\frac{R_{p}}{R_{\oplus}}=1.78~{}\Big{(}\frac{M_{p}}{M_{\oplus}}\Big{)}^{0.53}~{}\Big{(}\frac{F}{erg~{}s^{-1}~{}cm^{-2}}\Big{)}^{-0.03}~{}~{}~{}~{}~{}~{}\text{for}~{}~{}M_{p}<150~{}M_{\oplus}$ (4)
@@
where $F$ is the incident flux from the parent star and can be calculated by
@@
Table 1: Summary of the Adopted Planet Occurrence Frequencies, as a Function of Planet Mass and Period adopted from (Perryman et al., 2014; Wright et al., 2012)
Class Planet Mass $[M_{J}]$ Orbital Period $[day]$ Frequency
Earth 0.002-0.007 0.8 - 85 0.1840
Super Earth 0.007-0.018 0.8 - 145 0.2960
Small Neptune 0.018-0.033 0.8 - 245 0.3090
Large Neptune 0.033-0.077 0.8 - 418 0.0318
giant 0.077-0.3 0.8 - 418 0.0114
Restricted giant 0.1-0.3 418 - 10 yr 0.0388
giant 0.3-15 2 - 112 0.015
Giant Giant 0.3-15 112 - 10 yr 0.13
@@ 2.4 Estimation of Exoplanets Radio Emission
To simulate the detection process via Roman-SKA, a model is required to estimate the emission of exoplanets as a function of their characteristics and the features of the parent stars. In general, the flux received from a planet contains thermal radiation due to its intrinsic temperature, as well as the reflection radiation from the parent star and radio emission if it has a magnetic field due to the electron cyclotron maser instability (ECMI) (Melrose & Dulk, 1982; Dulk, 1985). The thermal and reflection radiation can be readily computed from the planet’s characteristics and its parent star (Bagheri et al., 2019). The ECMI radiation is characterized by strong circular polarization and high anisotropy. ECMI stands out as the most efficient mechanism for radio emission, superseding other mechanisms like beam-plasma instabilities. Consequently, it plays a dominant role in the generation of radio emissions. This dominance is exemplified in the brightness of auroral radio emissions from planets in the solar system, particularly Jupiter, which rival the intensity of solar radio bursts.
Based on observations of the magnetized Solar system planets, we know that all auroral radio emissions power are related to incident energy flux of the stellar wind (Zarka et al., 2001; Zarka, 2004, 2007; Zarka et al., 2018). In all cases in our solar system, about  1% of the electrons’ energy goes into radio waves (Zarka et al., 2018). Therefore, the average electron cyclotron maser instability can be linked to both the total power of the solar wind and the magnetic field strength of the planet, utilizing the RBL model (Zarka et al., 2001; Zarka, 2004, 2007). It can be expressed mathematically as:
@@
where $\nu$ is the emitted frequency, $e$ is the charge of the emitting particles (in this case, electrons), $B$ is the magnetic field strength, and $m$ is the mass of the emitting particles (electron mass in this case). This equation shows that the emitted frequency is directly proportional to the magnetic field strength of the planet. Therefore, observations of this emission type can provide an indirect means of measuring and classifying the planetary magnetic field. This quantification of the magnetic field gives us constraints on internal structure models and planetary rotation, informing us about the evolutionary history of the planet and host system (Hess et al., 2011).
The RBL model is frequently used to estimate the radio emission of extrasolar planets (Farrell et al., 1999; Zarka et al., 2001; Joseph et al., 2004; Zarka, 2007; Christensen, 2010). The simplicity of this model has some advantages, but it does not provide a complete picture of all processes involved as discussed in (Bagheri et al., 2024). However, the RBL model facilitates comparative analyses between the radio emissions of exoplanets and those observed within the solar system, thereby enhancing our contextual understanding of planetary magnetospheric dynamics. This capacity for comparative analysis underscores the utility of the RBL model in advancing our comprehension of exoplanetary environments and their potential habitability while also serving as a foundation for future investigations into the broader astrophysical implications of planetary radio emissions. Thus, in this paper, we follow the approach outlined by (Joseph et al., 2004) to estimate the radio flux density of exoplanets. As described by (Farrell et al., 1999), the radio power from an exoplanet is related to the incident power of the stellar winds. Following (Farrell et al., 1999), the Jovian decametric component is considered at least partly related to the solar wind kinetic energy input and is used as the base power level to give
@@
where $R_{\star}$ and $M_{\star}$ are the host star radius and mass (O’Gorman et al., 2018). Then, the radio flux density ($S_{\nu}$) emitted by an exoplanet can be estimated using the following equation
@@
where $R_{eq}$ and $R_{pole}$ are the equatorial and polar radii of the planet. In our simulation, for the Jovian planet with orbital distance $a>0.1$ au, we use the oblateness of Jupiter ($f=0.064$), and for rocky planets, we use the Earth’s value which is $f=0.00335$ (Barnes & Fortney, 2003).
@@
There are more than 200 moons in our solar system, many of which orbit gas giants like Jupiter and Saturn. Despite significant efforts, the successful detection of exomoons remains elusive (Narang et al., 2023). If an exomoon resides within the magnetosphere of its host planet, it can influence the planet’s radio emissions by injecting plasma that accelerates electrons within the magnetosphere. A notable example is Io, a moon characterized by intense volcanic activity within Jupiter’s magnetosphere. This activity generates a tenuous atmosphere around Io (Lopes et al., 2007), contributing ions to Jupiter’s magnetosphere and generating a plasma (Reich et al., 2009). These ions in motion accelerate electrons, ultimately leading to the formation of an ECMI along the field lines, resulting in radio emissions known as Io-DAM (2) (Crary, 1997; Mauk et al., 2001). Similar effects are observed with other Galilean moons and moons orbiting Saturn, such as Titan and Enceladus. Consequently, it is reasonable to anticipate moon-planet interactions in other stellar systems (Green et al., 2021). ECMI radio emissions from exomoon magnetosphere have not had a confirmed detection, but there are three candidate systems under study (Turner et al., 2021).
Figure 2: The radio emission of the planets in our solar system regarding frequency. Jupiter’s radio emission is classified into three categories: 1. radiation with frequencies less than 0.3 MHz is called the Jovian kilometric radiation or KOM; 2. radiation with frequencies in the interval of 0.3–3 MHz is called the hectometric radiation or HOM; and, 3. radiation in the range 3–40 MHz are referred to as the decametric radiation or DAM. The strongest part of decametric emission, which is related to Io–Jupiter interaction, is called Io-DAM. Figure adapted from (Cecconi, 2014)
To simulate the induced radio emission by exomoons, one should know the probability of having a moon such as Io. The results of N-body simulations in (Heller et al., 2014; Sasaki et al., 2010) show that 80% of Jovian planets have moons with masses $M_{m}/M_{p}~{}\approx 10^{-5}$. However, stellar tidal forces serve to slow down a planet’s rotation and in conjunction with tidal migration, contribute to removing moons; At short $a$, migration of moons is significant and occurs quickly, which can lead to orbital instabilities (Sucerquia et al., 2020). This process limits the lifetimes of larger moons in extrasolar planetary systems (Barnes & O’brien, 2002). Therefore, in our simulation, we assume 80% of Jovian planets with masses $M_{p}\geq 0.1~{}M_{J}$ with $a\geq 0.1\text{au}$ have one moon. So exoplanet with the semi-major axis less than 0.1 au do not have any moons, which in consistent with assuming they are tidally-locked.
@@
where $\nu_{c_{i}}$ is the frequency of radio emission induced by a moon, such as Io, $a_{m}$ is the same-major axis of the moon’s orbit, and $\nu_{c}$ can be obtained by (11). (17) is asymptotic, approaching a maximum value of $\nu_{c}$ as $a_{m}$ becomes very large (Noyola et al., 2016). Assuming the plasma density of the exomoon is equal to the Io’s plasma density, the optimal distance for the exomoon to have the maximum of the induced radio emission can be obtained by (Noyola et al., 2014)
@@
where $d$ is the distance of star-planet-moon system to Earth, $\rho_{m}$ is the moon’s plasma density, $R_{m}$ is the moon’s radius, and $\beta$ is the efficiency coefficient of converting Joule dissipation to radio waves (we assume $\beta\approx 1\%$ as is in the Jupiter–Io system (Zarka et al., 2001)). $B_{m}$ is the magnetic field affecting the exomoon, and so $B_{m}=B_{p}\times(R_{p}/a_{m})^{3}$, where $B_{p}$ is the magnetic field of the exoplanet at its poles. The plasma speed, $V_{0}$, is computed assuming it corotates with the planet’s magnetic field. Therefore, $V_{0}=\omega_{p}\times a_{m}-\sqrt{G~{}M_{p}/a_{m}}$, where $G$ is the gravitational constant, $M_{p}$is the planet’s mass, and $\omega_{p}$ is the planet’s corotation speed. Finally, $\rho_{c}$ is the critical density which is defined as $\rho_{c}=\mu_{0}^{-1}(B_{m}/V_{0})^{2}$ (Noyola et al., 2014). In our simulation, we assume IO’s plasma density and IO’s radius for $\rho_{m}$ and $R_{m}$. We also assume $\Omega=0.2$ (Noyola et al., 2014).

### v2

@@ Abstract
There are several methods for indirectly detecting exoplanets, such as transit, radial velocity, astrometry, and the conventional gravitational microlensing approach. These methods rely on observing the effects of exoplanets on the emission or motion of observed stars. All these techniques have focused on the optical or infrared domains. However, an alternative method for exoplanet detection via microlensing events involves planets orbiting the source star, creating a binary source system. In this study, we explore a novel approach to detecting and studying exoplanets exclusively through their radio emissions resulting from magnetospheric processes. We propose utilizing the Roman telescope as a survey observer to detect microlensing events. Subsequently, we investigate the potential for detecting planetary radio signals through follow-up observations of these microlensing events in the radio band using the SKA telescope. This method is viable due to the comparable radio emission levels of exoplanets and their parent stars, unlike optical and infrared emissions. We conduct a Monte Carlo simulation to replicate the observations by the Nancy Roman Telescope, followed by a follow-up observation in radio frequencies using the SKA telescope. We determine that approximately 1317 exoplanets exhibit detectable signals by the SKA telescope during the 7-season observations by the Nancy Roman Telescope. This result indicates that such a method cannot only facilitate the direct detection of exoplanets but also enable the measurement of their magnetic field strength through analysis of their radio emissions.
journal: Frontiers in Astronomy and Space Sciences
@@ 1 Introduction
A planetary magnetic field is one of the few observable quantities that contain information about the properties of the planetary interior since the internal dynamo produces it. Further, the extent to which exoplanets retain their atmospheres over geological time scales depends on whether they are exposed directly to stellar winds. Also, planetary magnetic fields could potentially serve as indicators of plate tectonics. Therefore, information about the magnetic fields of extrasolar planets is essential to assess their habitability. So far, our knowledge of planetary magnetic fields is limited to our solar system’s planets, which are not representative of the vast array of possible planets, such as hot Jupiters (Lammer et al., 2009; Lingam & Loeb, 2021; Cockell et al., 2016; Kane, 2021).
@@
In this work, we combine two different exoplanet detection methods to make the radio emission of exoplanets detectable. The novel idea behind this study is that during a microlensing event, light intensity will be magnified at all frequencies. Therefore, a microlensing event can serve as a natural light magnifier. The idea is to use an optical/infrared telescope to detect a microlensing event that would trigger follow-up observations by radio telescopes upon detection. If the source star has at least one exoplanet in microlensing events, we can directly measure the light emitted from the exoplanet (Figure 1). The magnified light of the exoplanet could be observable in the infrared wavelength as we have shown in (Bagheri et al., 2019). However, the radio band is the most desirable wavelength window to detect exoplanets (Vedantham et al., 2020; Viswanath et al., 2020; Foster et al., 2020) since the planetary radio emission, which includes the effects of exomoons, may dominate the signal from the parent star (Cauley et al., 2019; Zarka, 2004). This method will enable us to measure the magnetic fields of exoplanets and provide us with information about the existence of exomoons. To show this, we simulate the Nancy Grace Roman Space Telescope (previously named Wide Filed Infrared Space Telescope, WFIRST) as an infrared survey telescope to alert the Square Kilometre Array (SKA) radio telescope to the occurrence of microlensing events for follow-up observations.
@@ 2.1 Microlensing events
where $u$ is the impact parameter, defined as the closest approach of the source to the lens normalized by the Einstein radius (Gaudi, 2012). The lensing system may comprise either a single object (single-lens) or two objects (binary-lens), with the possibility of additional components (Daněk & Heyrovskỳ, 2019). The presence of two lenses introduces additional complexity to the gravitational lensing phenomenon, leading to more intricate and often asymmetric light curves. Source systems can also consist of one or more objects. The conventional technique utilized in gravitational microlensing for exoplanet detection relies on the premise that the lens star hosts an exoplanet, constituting a binary lens system. Alternatively, another microlensing approach involves a scenario where the planet orbits the source star, resulting in a binary source configuration (Bagheri et al., 2019) as illustrated in Figure 1. In this study, we simulate binary-source microlensing events by assuming both single-lens and binary-lens configurations. We consider the finite source effect for the source star while neglecting limb darkening.
@@ 2.3 Exoplanet Parameters and their Occurrence
In our simulations, we categorize the characteristic parameters of exoplanets into two primary groups: rocky planets and Jovian planets. According to the conventional planet formation theory, rocky planets are situated within the snow line of their parent star (Kennedy et al., 2006; Kennedy & Kenyon, 2008). The snowline of a star, also known as the frost line or ice line, refers to the distance from the star where volatile compounds, such as water, can condense into solid form due to the decrease in temperature. Beyond the snowline, the temperature drops low enough for these volatile compounds to freeze, forming icy grains or particles (Kennedy & Kenyon, 2008). This snowline can be expressed with the snowline of the Sun as
@@
The planetary mass range in our simulation is $0.002-10M_{J}$ (Schlaufman, 2018). For planets with masses $>100~{}M_{\oplus}$ orbiting host stars with masses greater than $0.6$ solar masses, we utilized the correlation between giant planet occurrence and host star mass and metallicity as outlined by (Johnson et al., 2010; Fulton et al., 2021). Following the occurrence dependencies discussed in (Fulton et al., 2021; Santerne et al., 2016, 2016; Wittenmyer et al., 2020; Winn & Petigura, 2024), we adopted their result of planet occurrence. We summarized the frequencies of various exoplanet types from our simulation in Table LABEL:tab:fvsMp.
@@
$\frac{R_{p}}{R_{\oplus}}=1.78~{}\Big{(}\frac{M_{p}}{M_{\oplus}}\Big{)}^{0.53}~{}\Big{(}\frac{F}{erg~{}s^{-1}~{}cm^{-2}}\Big{)}^{-0.03}~{}~{}~{}~{}~{}~{}\text{for}~{}~{}M_{p}<150~{}M_{\oplus}$
@@
where $F$ is the bolometric incident flux from the parent star and can be calculated by
@@
Table 1: Summary of the planet occurrence frequencies in our simulation, as a function of planet mass and period.
[…]
Earth 0.002-0.007 0.8 - $10^{4}$ 0.018
Super Earth 0.007-0.018 0.8 - $10^{4}$ 0.042
Small Neptune 0.018-0.033 0.8 $<$ 0.058
Large Neptune 0.033-0.077 10 $<$ 0.167
sub giant 0.077-0.3 10 $<$ 0.010
giant 0.3-5 10 $<$ 0.040
super giant 0.3-5 10 $<$ 0.49
Hot Jupiter 0.3-10 0.8 - 10 0.07
@@ 2.4 Estimation of Exoplanets Radio Emission
To simulate the detection process involving the Roman telescope for surveying microlensing events and subsequent follow-up observations with the SKA telescope, a model is necessary to estimate the emission of exoplanets based on their characteristics and the attributes of their parent stars. In general, the flux received from a planet contains thermal radiation due to its intrinsic temperature, as well as the reflection radiation from the parent star and radio emission if it has a magnetic field due to the electron cyclotron maser instability (ECMI) (Melrose & Dulk, 1982; Dulk, 1985). The thermal and reflection radiation can be readily computed from the planet’s characteristics and its parent star (Bagheri et al., 2019). The ECMI radiation is characterized by strong circular polarization and high anisotropy. ECMI stands out as the most efficient mechanism for radio emission, superseding other mechanisms like beam-plasma instabilities. Consequently, it plays a dominant role in the generation of radio emissions. This dominance is exemplified in the brightness of auroral radio emissions from planets in the solar system, particularly Jupiter, which rival the intensity of solar radio bursts.
Based on observations of the magnetized Solar system planets, we know that all auroral radio emissions power is related to incident energy flux of the stellar wind (Zarka et al., 2001; Zarka, 2004, 2007; Zarka et al., 2018). In all cases in our solar system, about  $2\times 10^{-3}$ of the electrons’ energy goes into radio waves (Zarka et al., 2018). Therefore, the average ECMI radiation can be linked to both the total power of the solar wind and the magnetic field strength of the planet, utilizing the RBL model (Zarka et al., 2001; Zarka, 2004, 2007). The gyrofrequency can be expressed mathematically as:
@@
where $\nu$ is the emitted pick frequency, $e$ is the charge of the emitting particles (in this case, electrons), $B$ is the magnetic field strength, and $m$ is the mass of the emitting particles (electron mass in this case). This equation shows that the emitted frequency is directly proportional to the magnetic field strength of the planet. Therefore, observations of this emission type can provide an indirect means of measuring and classifying the planetary magnetic field. This quantification of the magnetic field gives us constraints on internal structure models and planetary rotation, informing us about the evolutionary history of the planet and host system (Hess et al., 2011).
The RBL model is frequently used to estimate the radio emission of extrasolar planets (Farrell et al., 1999; Zarka et al., 2001; Joseph et al., 2004; Zarka, 2007; Christensen, 2010). The simplicity of this model has some advantages, but it does not provide a complete picture of all processes involved, in particular for close-in exoplanets as discussed in (Bagheri et al., 2024b) (for Examples of studies that extend beyond the RBL model for calculating radio emission (Vidotto & Donati, 2017; Li et al., 2023; Dong et al., 2018; Vidotto, 2017)). However, the RBL model facilitates comparative analyses between the radio emissions of exoplanets and those observed within the solar system, thereby enhancing our contextual understanding of planetary magnetospheric dynamics. This capacity for comparative analysis underscores the utility of the RBL model in advancing our comprehension of exoplanetary environments and their potential habitability while also serving as a foundation for future investigations into the broader astrophysical implications of planetary radio emissions. Thus, in this paper, we follow the approach outlined by (Joseph et al., 2004) to estimate the radio flux density of exoplanets. As described by (Farrell et al., 1999), the radio power from an exoplanet is related to the incident power of the stellar winds. Following (Farrell et al., 1999), the Jovian decametric component is considered at least partly related to the solar wind kinetic energy input and is used as the base power level to give
@@
where $R_{\star}$ and $M_{\star}$ are the host star radius and mass in solar radius and mass units (O’Gorman et al., 2018). Equation (8) serves as a mere approximation for low-mass stars, however, recent study (Mesquita & Vidotto, 2020) indicates that the terminal velocity of M dwarf stellar winds either aligns well with predictions from Parker wind models, in which stellar winds are driven by radiation pressure mechanism, or exceed these model estimates. As a result, employing this equation may yield conservative estimations of stellar winds terminal velocity and the overall count of planets with detectable signals. Then, the radio flux density ($S_{\nu}$) emitted by an exoplanet can be estimated using the following equation
@@
where $R_{eq}$ and $R_{pole}$ are the equatorial and polar radii of the planet. In our simulation, for the Jovian planet with orbital distance $a>0.1$ au, we use the oblateness of Jupiter ($f=0.064$), and for rocky planets, we use the Earth’s value which is $f=0.00335$ (Barnes & Fortney, 2003). In Our simulation we also assume no planets rotate faster than Jupiter.
@@
There are more than 200 moons in our solar system, many of which orbit gas giants like Jupiter and Saturn. Despite significant efforts, the successful detection of exomoons remains elusive (Narang et al., 2023). If an exomoon resides within the magnetosphere of its host planet, it can influence the planet’s radio emissions by injecting plasma that accelerates electrons within the magnetosphere. A notable example is Io, a moon characterized by intense volcanic activity within Jupiter’s magnetosphere. This activity generates a tenuous atmosphere around Io (Lopes et al., 2007), contributing ions to Jupiter’s magnetosphere and generating a plasma (Reich et al., 2009). These ions in motion accelerate electrons, ultimately leading to the formation of an ECMI along the field lines, resulting in radio emissions known as Io-DAM (see Figure 1-a in (Cecconi, 2014)) (Crary, 1997; Mauk et al., 2001). Similar effects are observed with other Galilean moons and moons orbiting Saturn, such as Titan and Enceladus. Consequently, it is reasonable to anticipate moon-planet interactions in other stellar systems (Green et al., 2021). ECMI radio emissions from exomoon magnetosphere have not had a confirmed detection, but there are three candidate systems under study (Turner et al., 2021).
To simulate the induced radio emission by exomoons, one should know the probability of having a moon such as Io. The results of N-body simulations in (Heller et al., 2014; Sasaki et al., 2010) show that 80% of Jovian planets have moons with masses $M_{m}/M_{p}~{}\approx 10^{-5}$. However, stellar tidal forces serve to slow down a planet’s rotation and in conjunction with tidal migration, contribute to removing moons; At short $a$, migration of moons is significant and occurs quickly, which can lead to orbital instabilities (Sucerquia et al., 2020). This process limits the lifetimes of larger moons in extrasolar planetary systems (Barnes & O’brien, 2002). Therefore, in our simulation, we assume 80% of Jovian planets with masses $M_{p}\geq 0.1~{}M_{J}$ with $a\geq 0.1\text{au}$ have one moon. So exoplanets with the semi-major axis of less than 0.1 au do not have any moons, which is consistent with assuming they are tidally-locked.
@@
where $\nu_{c_{i}}$ is the frequency of radio emission induced by a moon, such as Io, $a_{m}$ is the same-major axis of the moon’s orbit, and $\nu_{c}$ can be obtained by (10). Equation (16) is asymptotic, approaching a maximum value of $\nu_{c}$ as $a_{m}$ becomes very large (Noyola et al., 2016). Assuming the plasma density of the exomoon is equal to the Io’s plasma density, the optimal distance for the exomoon to have the maximum of the induced radio emission can be obtained by (Noyola et al., 2014)
@@
where $d$ is the distance of star-planet-moon system to Earth, $\rho_{m}$ is the moon’s plasma density, $R_{m}$ is the moon’s radius, and $\beta$ is the efficiency coefficient of converting Joule dissipation to radio waves (we assume $\beta\approx 1\%$ as is in the Jupiter–Io system (Zarka et al., 2001)). $B_{m}$ is the magnetic field affecting the exomoon, and so $B_{m}=B_{p}\times(R_{p}/a_{m})^{3}$, where $B_{p}$ is the magnetic field of the exoplanet at its poles. The plasma speed, $V_{0}$, is computed assuming it corotates with the planet’s magnetic field. Therefore, $V_{0}=\omega_{p}\times a_{m}-\sqrt{G~{}M_{p}/a_{m}}$, where $G$ is the gravitational constant, $M_{p}$is the planet’s mass, and $\omega_{p}$ is the planet’s corotation speed. Finally, $\rho_{c}$ is the critical density which is defined as $\rho_{c}=\mu_{0}^{-1}(B_{m}/V_{0})^{2}$ (Noyola et al., 2014). In our simulation, we assume IO’s plasma density and IO’s radius for $\rho_{m}$ and $R_{m}$. We also assume $\Omega=0.2$ (Noyola et al., 2014). In our simulation, the induced radio emission by an Io-like moon accounts for a maximum of 0.2% of the total radiation emitted by the exoplanets.